# OccFusion: A Straightforward and Effective Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### 13 Mar 2024 v2

@@ Abstract
This paper introduces OccFusion, a straightforward and efficient sensor fusion framework for predicting 3D occupancy. A comprehensive understanding of 3D scenes is crucial in autonomous driving, and recent models for 3D semantic occupancy prediction have successfully addressed the challenge of describing real-world objects with varied shapes and classes. However, existing methods for 3D occupancy prediction heavily rely on surround-view camera images, making them susceptible to changes in lighting and weather conditions. By integrating features from additional sensors, such as lidar and surround view radars, our framework enhances the accuracy and robustness of occupancy prediction, resulting in top-tier performance on the nuScenes benchmark. Furthermore, extensive experiments conducted on the nuScenes dataset, including challenging night and rainy scenarios, confirm the superior performance of our sensor fusion strategy across various perception ranges. The code for this framework will be made available at https://github.com/DanielMing123/OCCFusion.
@@
autonomous driving, 3D semantic occupancy prediction, environment perception
@@ I Introduction
Understanding and modelling the three-dimensional (3D) world of surrounding scenes is essential for autonomous driving to prevent collisions and facilitate local planning. As technology advances, the introduction of 3D occupancy representation has successfully addressed the limitations of traditional 3D object detection networks, particularly for detecting irregular objects and space-occupied status prediction. This advancement further enhances the capability of smart vehicles to model the 3D world. However, the currently proposed 3D occupancy prediction models [1, 2, 3, 4, 5, 6, 7, 6, 8, 9, 10, 11, 12, 13, 14], primarily focus on vision-based approaches (Figure 1, upper). While surround-view cameras are cost-effective, their perception capabilities are highly susceptible to various weather conditions, such as sunny, rainy, foggy, and different illumination conditions, like daytime and nighttime. These factors cause the model to perform inconsistently in these scenarios, thereby posing potential safety risks.
@@
Besides the surround-view cameras, modern smart vehicles are also equipped with advanced lidar and surround-view millimetre wave radars. Lidar excels at capturing the geometric shape of objects and accurately measuring depth. Moreover, it is resistant to changes in illumination and performs reliably under various weather conditions, except for heavy rain and fog. In contrast, surround-view millimetre wave radars are cost-effective and exceptionally robust against weather conditions and changes in illumination. However, they yield only sparse features, which are often noisy. Each sensor possesses its advantages and disadvantages. We argue that combining information from all three sensors (Figure 1, bottom) can empower smart vehicles’ 3D occupancy prediction model to achieve superior comprehension and modelling of the 3D world. Additionally, this integration improves the system’s robustness in the face of varying lighting and weather conditions.
@@
Figure 1: The figure above demonstrates the pipeline of two approaches: The purely vision-centric approach (top) and the multi-sensor fusion approach (bottom). We conduct 3D semantic occupancy prediction by doing feature fusion with respect to three modality feature volumes.
@@
To predict 3D occupancy accurately and efficiently, we propose a framework called OccFusion. This framework merged features from surround-view cameras, surround radars, and 360 degree lidar through proposed dynamic fusion 3D/2D modules. There are three sensor fusion strategies, namely, Camera + Radar, Camera + Lidar, and Camera + Lidar + Radar. To evaluate the effectiveness of our sensor fusion strategies, we conduct experiments on the nuScenes dataset [15], utilizing ground truth labels provided by the SurroundOcc [7] work. Additionally, we manually select rainy and nighttime scenarios from the nuScenes validation set to create two challenging subsets and examine the performance of different sensor fusion strategies in these scenarios. Finally, we test how the performance of different sensor fusion strategies varies across different perception ranges under different scenarios.
The main contributions of this paper are summarized below.
@@
  • A multi-sensor fusion framework is proposed to integrate camera, lidar, and radar information for the task of 3D semantic occupancy prediction.
@@
  • We conducted thorough ablation studies to assess the performance gains achieved by different sensor combinations under challenging lighting and weather conditions, such as night and rainy scenarios.
@@ II Related Work
II-A Camera-only environment perception
@@
In recent years, surround-view camera-based environment perception algorithms have received significant attention in the autonomous driving domain due to their cost-effectiveness and versatility. The BEV feature-based algorithms [16, 2, 17, 4, 3, 18, 19] successfully merged all surround-view camera information to conduct 3D object detection tasks. By lifting the BEV feature into 3D feature volume, algorithms [2, 10] are capable of doing 3D semantic occupancy tasks. There are two main approaches to view transformation in these algorithms: the classic Lift-Splat-Shoot (LSS) approach and the transformer-based approach. The LSS-based approach [12, 20] relies on depth estimation to generate a pseudo-3D point cloud, followed by voxel-pooling to create the final 3D feature volume. On the other hand, the transformer-based approach [5, 11, 6, 13, 8, 9, 7] uses sampling points to aggregate visual features from feature maps and places these features directly at specific 3D positions in the world to form the final 3D feature volume. Both approaches either explicitly estimate the depth or implicitly encode depth information in visual features. Nonetheless, it is well-known that monocular cameras are inadequate for accurate depth estimation. While they can capture the relative depth position of an object, they cannot provide precise depth information. Hence, a more reliable reference for depth information is required. This could involve incorporating lidar information into the model to enhance depth estimation or using lidar information to supervise depth estimation, as in the BEVDet series approaches.
II-B Lidar-only environment perception
@@
Lidar-based algorithms [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] for environment perception have shown promising performance in various perception tasks. Leveraging its capability for accurate depth estimation, lidar excels in capturing the geometric shape and 3D location of objects. By converting the 3D point cloud into Euclidean feature spaces, such as 3D voxel grids [32] or feature pillars [33], lidar-based methods can achieve highly precise 3D object detection results. In recent years, researchers have extended lidar’s 3D point cloud features into 3D semantic occupancy prediction tasks [34, 35]. However, the density of the lidar-generated 3D point cloud strongly influences the final perception performance of the model, and its lack of semantic information results in inaccurate object class recognition. Hence, the auxiliary information is needed to provide comprehensive semantic information guidance, which leads to our work fusing the lidar data with camera data to enhance the performance of 3D semantic occupancy prediction.
II-C Camera-Lidar fusion environment perception
@@
Due to the inherent advantages and disadvantages of individual sensors, recent research has focused on sensor fusion techniques [36, 37, 38, 39, 40, 41] to overcome these limitations and enhance the overall environment perception capability of the models. The representative BEVFusion [42, 43] algorithms fuse lidar and surround-view cameras by encoding the features of each modality into BEV features and performing feature fusion. This approach addresses the reflection issues encountered by lidar in rainy and foggy scenarios, which often result in false detections and missed detections. It also resolves the poor depth estimation problem of monocular cameras, enabling the model to generate relatively accurate detection results at longer distances. The SparseFusion [44] further refines the inner structure of the feature fusion module, greatly improving the model’s inference speed. However, it is regrettable that currently there are no camera-lidar fusion algorithms specifically designed for the 3D semantic occupancy prediction task; the majority of existing algorithms primarily serve for 3D object detection. Therefore, there is a compelling need for extensive research on camera-lidar fusion techniques for 3D semantic occupancy prediction.
II-D Camera-Radar fusion environment perception
@@
Various studies [45, 46, 47] have been conducted on the fusion of cameras and radar for environmental perception due to radar’s cost-effectiveness and ability to detect distant objects. For instance, the work carried out in [48] demonstrates that information about velocity obtained from radar sensors can enhance detection performance. Additionally, a study by [49] suggests that integrating radar features with visual features can result in a performance gain of approximately 12% under the nuScene Detection Score (NDS) metric. Furthermore, another study in [50] found that radar sensor readings exhibit robustness in noisy conditions, and integrating radar information can improve model performance in challenging scenarios. Though various algorithms have been developed to achieve camera-radar fusion, the majority of these algorithms focus on 3D object detection, object tracking, and object future trajectory prediction tasks. There is currently no camera-radar fusion algorithm available specifically for 3D semantic occupancy prediction tasks. Moreover, the task of 3D semantic occupancy prediction requires dense features, whereas radar provides sparse features. Consequently, it is both intriguing and valuable to investigate the impact on model performance when merging these sparse radar features with camera and lidar data. To the best of our knowledge, our study is the first to examine the influence of fused radar information on a 3D semantic occupancy prediction task.
II-E Camera-Lidar-Radar fusion environment perception
@@
Due to the complementary property of multi-sensor fusion, people in this domain also investigate the camera-lidar-radar fusion strategy and its performance in environment perception. In CLR-BNN [51], the authors employ a Bayesian neural network for camera-lidar-radar sensor fusion, yielding improved 2D multi-object detection results in terms of mAP. In Futr3D [52], the sensor fusion is further explored through the incorporation of transformers, using sparse 3D points as queries to aggregate features from the three sensors for the 3D object detection task, the query-form significantly improved feature interactions and aggregation efficiency across three sensors. In SimpleBEV [53], authors process data from all three sensors into Bird’s Eye View (BEV) features, fusing these features based on three BEV representations to perform the 3D object detection task; in their research, they found radar data provide a substantial boost to performance. Previous research has extensively examined the characteristics of three sensor fusion methods for environmental perception. However, these works have primarily concentrated on 2D or 3D multi-object detection tasks, neglecting investigations into the 3D semantic occupancy prediction task. Therefore, there is a need to investigate the performance of the camera-lidar-radar sensor fusion strategy in the context of 3D semantic occupancy prediction tasks.
@@ III OccFusion
Figure 2: Overall architecture of OccFusion. Firstly, the surround-view images were inputted into the 2D backbone to extract multiple-scale features. Subsequently, a view transformation is conducted at each scale to obtain the global BEV feature and the local 3D feature volume at each level. The 3D point cloud generated by the lidar and surround-view radars is also inputted into the 3D backbone to generate multi-scale local 3D feature volumes and global BEV features. The dynamic fusion 3D/2D modules at each level fuse features from the cameras and lidar/radar. Following this, the merged global BEV feature and local 3D feature volume at each level are fed into the global-local attention fusion to generate the final 3D volume at each scale. Finally, the 3D volume at each level is upsampled, and the skip connection is performed while adopting a multi-scale supervision mechanism.
@@ III-A Problem Statement
where $F$ represents the fusion framework that integrates multi-sensor information for 3D occupancy prediction. The final 3D occupancy prediction result is represented by $Occ\in R^{X\times Y\times Z}$, where each grid is assigned a semantic property and has a value ranging from 0 to 16. In our case, a class value of 0 corresponds to an empty grid.
@@ III-B Overall Architecture
Figure 2 exhibits the overall architecture of our proposed framework. Initially, given surround-view images, dense 3D point clouds $P^{Dense}$ from lidar and sparse 3D point clouds $P^{Sparse}$ from surround-view radars, we apply a 2D backbone (e.g. ResNet101-DCN) to extract total $L$ scale features $M=\left\{\left\{M_{n}^{l}\right\}_{n=1}^{N}\in R^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L}$ from images, followed by view transformation to obtain the global BEV feature $F_{global}^{Cam\_l}\in R^{C_{l}\times X_{l}\times Y_{l}}$ and local 3D feature volume $F_{local}^{Cam\_l}\in R^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$ at each scale. Meanwhile, a 3D backbone (e.g. VoxelNet) is also applied to dense and sparse 3D point clouds to generate multi-scale global BEV features $F_{global}^{Rad\_l}\in R^{C_{l}\times X_{l}\times Y_{l}}$, $F_{global}^{Lid\_l}\in R^{C_{l}\times X_{l}\times Y_{l}}$ and local 3D feature volumes $F_{local}^{Rad\_l}\in R^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$, $F_{local}^{Lid\_l}\in R^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$, respectively. Following that, the $F_{global}^{Cam\_l}$, $F_{global}^{Rad\_l}$ and $F_{global}^{Lid\_l}$ at each level are fed into the dynamic fusion 2D module to obtain the merged global BEV feature $F_{global}^{Merged\_l}$. Simultaneously, the $F_{local}^{Cam\_l}$, $F_{local}^{Rad\_l}$ and $F_{local}^{Lid\_l}$ at each level are also fed into the dynamic fusion 3D module to obtain the merged local 3D feature volume $F_{local}^{Merged\_l}$. Subsequently, the global-local attention fusion module proposed in [14] is used at each level to further merge $F_{local}^{Merged\_l}$ and $F_{global}^{Merged\_l}$, resulting in the final 3D volume at each level. Moreover, a skip-connection structure is implemented between each level to refine the features in a coarse-to-fine manner, and multi-scale supervision is applied to improve the model’s performance.

### 09 May 2024 v4

# OccFusion: Multi-Sensor Fusion Framework for 3D Semantic Occupancy Prediction

@@ Abstract
A comprehensive understanding of 3D scenes is crucial in autonomous vehicles (AVs), and recent models for 3D semantic occupancy prediction have successfully addressed the challenge of describing real-world objects with varied shapes and classes. However, existing methods for 3D occupancy prediction heavily rely on surround-view camera images, making them susceptible to changes in lighting and weather conditions. This paper introduces OccFusion, a novel sensor fusion framework for predicting 3D occupancy. By integrating features from additional sensors, such as lidar and surround view radars, our framework enhances the accuracy and robustness of occupancy prediction, resulting in top-tier performance on the nuScenes benchmark. Furthermore, extensive experiments conducted on the nuScenes and semanticKITTI dataset, including challenging night and rainy scenarios, confirm the superior performance of our sensor fusion strategy across various perception ranges. The code for this framework will be made available at https://github.com/DanielMing123/OccFusion.
@@
autonomous vehicles, 3D semantic occupancy prediction, environment perception
@@ I Introduction
Understanding and modelling the three-dimensional (3D) world is essential for autonomous vehicles (AVs) to navigate safely, preventing collisions and facilitating local planning. As technology advances, the introduction of 3D occupancy representation has successfully addressed the limitations of traditional 3D object detection networks, particularly for detecting irregular objects and space-occupied status prediction. This advancement further enhances the capability of AVs to model the 3D world. However, the currently proposed 3D occupancy prediction models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], primarily focus on vision-based approaches (Figure 1, upper). While surround-view cameras are cost-effective, their perception capabilities are highly susceptible to poor weather conditions like rain or fog, and illumination conditions such as those experienced at night. These factors cause the model to perform inconsistently in these scenarios, posing potential safety risks.
@@
Besides the surround-view cameras, AVs are often equipped with lidars and surround-view millimetre wave radars. Lidar excels at capturing the geometric shape of objects and accurately measuring depth. Moreover, it is robust to illumination changes and performs reliably under various weather conditions, except for heavy rain and fog. In contrast, surround-view millimetre wave radars are cost-effective and exceptionally robust against weather conditions and changes in illumination. However, they yield only sparse features, which are often noisy. Each sensor has its advantages and disadvantages. Combining information from all three sensors (Figure 1, bottom) can potentially enhance AVs’ 3D occupancy prediction model to achieve superior accuracy when modelling the 3D world. Additionally, this integration improves the system’s robustness in the face of varying lighting and weather conditions.
@@
Figure 1: Pipeline for two approaches: purely vision-centric approach (top) and multi-sensor fusion approach (bottom). We conduct 3D semantic occupancy prediction by doing feature fusion with respect to three modality feature volumes.
@@
To predict 3D occupancy accurately and efficiently, we propose a framework called OccFusion. This framework merged features from surround-view cameras, surround radars, and 360-degree lidar through dynamic fusion 3D/2D modules. There are three sensor fusion strategies: Camera + Radar, Camera + Lidar, and Camera + Lidar + Radar. To evaluate the accuracy of our sensor fusion strategies, we conduct experiments on the nuScenes dataset [15], utilizing ground truth labels provided by the SurroundOcc [7] and Occ3D [6]. Additionally, we manually select rainy and nighttime scenarios from the nuScenes validation set to create two challenging subsets and examine the performance of different sensor fusion strategies in these scenarios. Finally, we test how the performance of different sensor fusion strategies varies across different perception ranges under different scenarios.
The main contributions of this paper are summarized as below.
@@
  • A multi-sensor fusion framework is proposed to integrate camera, lidar, and radar information for the 3D semantic occupancy prediction task.
@@
  • We conducted thorough ablation studies to assess the performance gains of different sensor combinations under challenging lighting and weather conditions, such as night and rainy scenarios.
@@ II Related Work
This section presents the most recent research findings on various sensor fusion algorithms used for environmental perception in autonomous driving contexts.
II-A Camera-only-based environment perception
@@
In recent years, surround-view camera-based environment perception algorithms have received significant attention in the AVs domain due to their cost-effectiveness and versatility. The bird’s-eye-view (BEV) feature-based algorithms [16, 2, 17, 4, 3, 18, 19] successfully merged all surround-view camera information to conduct 3D object detection tasks. By lifting the BEV feature into 3D feature volume, algorithms [2, 10] are capable of doing 3D semantic occupancy tasks. Two main approaches to view transformation in these algorithms are the classic Lift-Splat-Shoot (LSS) and transformer-based approaches. The LSS-based approach [12, 20] relies on depth estimation to generate a pseudo-3D point cloud, followed by voxel-pooling to create the final 3D feature volume. On the other hand, the transformer-based approach [5, 11, 6, 13, 8, 9, 7] uses sampling points to aggregate visual features from feature maps and places these features directly at specific 3D positions in the world to form the final 3D feature volume. Both approaches explicitly estimate the depth or implicitly encode depth information in visual features. Nonetheless, it is well-known that monocular cameras are inadequate for accurate depth estimation. While they can capture the relative depth position of an object, they cannot provide precise depth information. Hence, a more reliable reference for depth information is required. This could involve incorporating lidar information into the model to enhance depth estimation or using lidar information to supervise depth estimation, as in the BEVDet series approaches.
II-B Lidar-only-based environment perception
@@
Lidar-only-based algorithms [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] for environment perception have shown promising performance in various perception tasks. Leveraging its capability for accurate depth estimation, lidar excels in capturing the geometric shape and 3D location of objects. By converting the 3D point cloud into Euclidean feature spaces, such as 3D voxel grids [32] or feature pillars [33], lidar-based methods can achieve highly precise 3D object detection results. In recent years, researchers have extended lidar’s 3D point cloud features into 3D semantic occupancy prediction tasks [34, 35]. However, the density of the lidar-generated 3D point cloud strongly influences the final perception performance of the model, and its lack of semantic information results in inaccurate object class recognition. Hence, the auxiliary information is needed to provide comprehensive semantic information guidance, which leads to our work fusing the lidar data with camera data to enhance the performance of 3D semantic occupancy prediction.
II-C Camera-Lidar fusion-based environment perception
@@
Due to the inherent advantages and disadvantages of individual sensors, recent research has focused on sensor fusion techniques [36, 37, 38, 39, 40, 41] to overcome these limitations and enhance the overall environment perception capability of the models. The representative BEVFusion [42, 43] algorithms fuse lidar and surround-view cameras by encoding the features of each modality into BEV features and performing feature fusion. This approach addresses the reflection issues lidar encounters in rainy and foggy scenarios, often resulting in false and missed detections. It also resolves monocular cameras’ poor depth estimation problem, enabling the model to generate relatively accurate detection results at longer distances. The SparseFusion [44] further refines the inner structure of the feature fusion module, greatly improving the model’s inference speed. However, currently, most existing algorithms primarily serve for 3D object detection. Therefore, there is a compelling need for extensive research on camera-lidar fusion techniques for 3D semantic occupancy prediction.
II-D Camera-Radar fusion-based environment perception
@@
Various studies [45, 46, 47] have been conducted on the fusion of cameras and radar for environmental perception due to radar’s cost-effectiveness and ability to detect distant objects. For instance, the work in [48] demonstrates that information about velocity obtained from radar sensors can enhance detection performance. Additionally, a study by [49] suggests that integrating radar features with visual features can result in a performance gain of approximately 12% under the nuScene Detection Score (NDS) metric. Furthermore, another study in [50] found that radar sensor readings exhibit robustness in noisy conditions, and integrating radar information can improve model performance in challenging scenarios. Though various algorithms have been developed to achieve camera-radar fusion, most focus on 3D object detection, object tracking, and object future trajectory prediction tasks. No camera-radar fusion algorithm is available specifically for 3D semantic occupancy prediction tasks. Moreover, this task requires dense features, whereas radar provides sparse features. Consequently, we investigate the impact on model performance when merging these sparse radar features with camera and lidar data. To the best of our knowledge, our study is the first to examine the influence of fused radar information on a 3D semantic occupancy prediction task.
II-E Camera-Lidar-Radar fusion-based environment perception
@@
Due to the complementary property of multi-sensor fusion, people in this domain also investigate the camera-lidar-radar fusion strategy and its performance in environment perception. In CLR-BNN [51], the authors employ a Bayesian neural network for camera-lidar-radar sensor fusion, yielding improved 2D multi-object detection results in terms of mAP. In Futr3D [52], the sensor fusion is further explored through the incorporation of transformers, using sparse 3D points as queries to aggregate features from the three sensors for the 3D object detection task, the query-form significantly improved feature interactions and aggregation efficiency across three sensors. In SimpleBEV [53], authors process data from all three sensors into BEV features, fusing these features based on three BEV representations to perform the 3D object detection task; in their research, they found radar data provide a substantial boost to performance. Previous research has extensively examined the characteristics of three sensor fusion methods for environmental perception. However, these works have primarily concentrated on 2D or 3D multi-object detection tasks, neglecting investigations into the 3D semantic occupancy prediction task. Therefore, there is a need to investigate the performance of the camera-lidar-radar sensor fusion strategy in the context of 3D semantic occupancy prediction tasks.
@@ III OccFusion
Figure 2: Overall architecture of OccFusion. Firstly, the surround-view images were inputted into the 2D backbone to extract multiple-scale features. Subsequently, each scale’s view transformation is conducted to obtain each level’s global BEV feature and the local 3D feature volume. The 3D point cloud generated by the lidar and surround-view radars is also inputted into the 3D backbone to generate multi-scale local 3D feature volumes and global BEV features. The dynamic fusion 3D/2D modules at each level fuse features from the cameras and lidar/radar. Following this, each level’s merged global BEV feature and local 3D feature volume are fed into the global-local attention fusion to generate the final 3D volume at each scale. Finally, the 3D volume at each level is upsampled, and the skip connection is performed while adopting a multi-scale supervision mechanism.
@@ III-A Problem Statement
where $F$ represents the fusion framework integrating multi-sensor information for 3D occupancy prediction. The final 3D occupancy prediction result is represented by $Occ\in R^{X\times Y\times Z}$, where each grid is assigned a semantic property ranging from 0 to 17. In our case, a class value of 0 corresponds to an empty grid.
@@ III-B Overall Architecture
Figure 2 exhibits the overall architecture of our proposed framework. Initially, given surround-view images, dense 3D point clouds $P^{Dense}$ from lidar and sparse 3D point clouds $P^{Sparse}$ from surround-view radars, we apply a 2D backbone (e.g. ResNet101-DCN) to extract total $L$ scale features $M=\left\{\left\{M_{n}^{l}\right\}_{n=1}^{N}\in R^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L}$ from images, followed by view transformation to obtain the global BEV feature $F_{global}^{Cam\_l}\in R^{C_{l}\times X_{l}\times Y_{l}}$ and local 3D feature volume $F_{local}^{Cam\_l}\in R^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$ at each scale. Meanwhile, a 3D backbone (e.g. VoxelNet) is also applied to dense and sparse 3D point clouds to generate multi-scale global BEV features $F_{global}^{Rad\_l}\in R^{C_{l}\times X_{l}\times Y_{l}}$, $F_{global}^{Lid\_l}\in R^{C_{l}\times X_{l}\times Y_{l}}$ and local 3D feature volumes $F_{local}^{Rad\_l}\in R^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$, $F_{local}^{Lid\_l}\in R^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$, respectively. Following that, the $F_{global}^{Cam\_l}$, $F_{global}^{Rad\_l}$ and $F_{global}^{Lid\_l}$ at each level are fed into the dynamic fusion 2D module to obtain the merged global BEV feature $F_{global}^{Merged\_l}$. Simultaneously, the $F_{local}^{Cam\_l}$, $F_{local}^{Rad\_l}$ and $F_{local}^{Lid\_l}$ at each level are also fed into the dynamic fusion 3D module to obtain the merged local 3D feature volume $F_{local}^{Merged\_l}$. Subsequently, the global-local attention fusion module proposed in [14] is used at each level to merge further $F_{local}^{Merged\_l}$ and $F_{global}^{Merged\_l}$, resulting in the final 3D volume at each level. Moreover, a skip-connection structure is implemented between each level to refine the features in a coarse-to-fine manner, and multi-scale supervision is applied to improve the model’s performance.